OpenStudy (prince1342):

The equation of line CD is y = 3x − 3. Write an equation of a line perpendicular to line CD in slope-intercept form that contains point (3, 1).

OpenStudy (prince1342):

y = 3x + 0 y = 3x − 8 y = negative 1 over 3x + 2 y = negative 1 over 3x + 0

OpenStudy (prince1342):

I'm so lost :(

OpenStudy (prince1342):

@Vijeya3

OpenStudy (prince1342):

@gottennis121

OpenStudy (gottennis121):

Gee, I just became brain dead due to telling myself I'm done with school for the day. I know the perfect person to ask though. Please welcome, @mathmath333 !! Ok, you get the point. I'm sorry, I probably know this when I'm in the zone.

OpenStudy (prince1342):

It's ok, thanks for referring me to someone, i really appreciate it!

OpenStudy (gottennis121):

No prob. I wish my hamster wheel was working, but a bit screechy. xD

OpenStudy (prince1342):

Hahah

OpenStudy (faiqraees):

are you here?

OpenStudy (prince1342):

Yep

OpenStudy (faiqraees):

okay first you have to work out the gradient of the second line We know its perpendicular to the first line which means their gradients follow the property \[m1 *m2 = -1\] Using this property can you calculate the gradient of the next line?

OpenStudy (prince1342):

Do I plug in the coordinates (3,1)

OpenStudy (faiqraees):

no you have to plug in the gradient of the first line m means gradient m1 means gradient of first line m2 means gradient of second line

OpenStudy (prince1342):

How do i find the gradient? Haven't heard that word used in math

OpenStudy (faiqraees):

gradient also means slope of a line

OpenStudy (faiqraees):

now can you figure it out?

OpenStudy (anonymous):

well attach the doc on plz so i can see original problem

OpenStudy (prince1342):

Hmmm so I have to get the slope of the coordinates first?

OpenStudy (prince1342):

@Sohal Here you go

OpenStudy (faiqraees):

slope of the line is the constant attached to x

OpenStudy (prince1342):

Oh okay, so 3?

OpenStudy (faiqraees):

yes

OpenStudy (prince1342):

Making the gradient of the first line 3?

OpenStudy (faiqraees):

plug it in the equation i gave you to find out the slope of the other line

OpenStudy (prince1342):

it would give me 3 again

OpenStudy (faiqraees):

no it would give you 3 * m2 =-1 m2 = -1/3

OpenStudy (prince1342):

Ohhh so because we dont have m2 we reverse the equation by divding 3 by -1

OpenStudy (prince1342):

Giving us -3

OpenStudy (faiqraees):

now you have the slope (or gradient or change in y/ change in x) use it in the equation y = mx +c where m is the m2 and c is the constant you will find by plugging in the x = 3 and y =1

OpenStudy (prince1342):

Hmm im a little lost again sorry

OpenStudy (faiqraees):

y = mx +c is the equation you will use now plug the value of m2 in its place. WHats the equation now?

OpenStudy (prince1342):

y=-3+c

OpenStudy (faiqraees):

y = -1/3 x +c

OpenStudy (faiqraees):

\[y = -1/3 *x +c\]

OpenStudy (faiqraees):

now input the value x =3 and y =1 and find out the value of c

OpenStudy (prince1342):

Hm okay, ill try that out

OpenStudy (prince1342):

1=-1/3 3 + c

OpenStudy (prince1342):

So what would that make C?

OpenStudy (faiqraees):

1 = -1 +c c =2 y =( -1/3 )x +2

OpenStudy (prince1342):

Ohhh okay I get it, thanks!!

OpenStudy (faiqraees):

good